OpenStudy (anonymous):

if k belongs to real no. and middle term of (k/2+2)^8 is 1120 . find alue of k.

OpenStudy (cwrw238):

the middle term is the 5th. the (r+1)th term of (a + x)^2 = nCr a^(n-r) x^ r here r = 4, a = (k/2) and x = 2 , n = 8

OpenStudy (cwrw238):

* typing error the (r+1)th term of (a + x)^n = nCr a^(n-r) x^ r